# Receptogenesis in a Vascularized Robotic Embodiment

### 10 Mar 2026 v1

@@ Abstract
Equipping robotic systems with the capacity to generate ex novo hardware during operation extends control of physical adaptability. Unlike modular systems that rely on discrete component integration pre- or post-deployment, we envision the possibility that physical adaptation and development emerge from dynamic material restructuring to shape the body’s intrinsic functions. Drawing inspiration from circulatory systems that redistribute mass and function in biological organisms, we utilize fluidics to restructure the material interface, a capability currently unpaired in robotics. Here, we realize this synthetic growth capability through a vascularized robotic composite designed for programmable material synthesis, demonstrated via receptogenesis - the on-demand construction of sensors from internal fluid reserves based on environmental cues. By coordinating the fluidic transport of precursors with external localized UV irradiation, we drive an in situ photopolymerization that chemically reconstructs the vasculature from the inside out. This reaction converts precursors with photolatent initiator into a solid dispersion of UV-sensitive polypyrrole, establishing a sensing modality validated by a characteristic decrease in electrical impedance. The newly synthesized sensor closed a control loop to regulate wing flapping in a moth-inspired robotic demonstrator. This physical update increased the robot’s capability in real time. This work establishes a materials-based framework for constitutive evolution, enabling robots to physically grow the hardware needed to support emerging behaviors in a complex environment; for example, suggesting a pathway toward autonomous systems capable of generating specialized features, such as neurovascular systems in situated robotics.
@@ 1 Introduction
In unstructured environments, critical physical signals are not static features that can be precisely registered by a dedicated sensor, but rather emergent phenomena arising from the robot’s unique interactions with the environment [27]. This unpredictability exposes the insufficiency of static hardware, necessitating systems capable of constitutive evolution to physically instantiate the features needed to transduce these unique, situation-specific signals [1]. We need more focus on situated robots [41] that are deeply intertwined with its immediate surroundings and leverage the environment itself to guide its behaviour and expand its physical capabilities. Current robotics lacks morphogenetic mechanisms that couple environmental information with physical updates for expanding functionality.
In previous works, physical adaptability during deployment has been realised through compliance/deformation [33], configuration change (rearrangement) [37, 7, 28], growth [34], or software solutions[22]. Multi-substrate environments could be traversed by increasing locomotion plasticity with multi-purpose appendages that could be configured as thrusters, wheels, and legs [37]. Similarly, vastly different environments, such as aquatic and terrestrial, could be transversed with shape-morphing limbs [7] or body-wide morphological reconfiguration [28]. Remapping and fusing sensory feed according to situational awareness is a sign of robustness in current robots, e.g., by repurposing navigational accelerometers for gravimetry in the Curiosity rover [22]. A form of temporal adaptability is self-healing, e.g., recovering from damage by harnessing embedded polymer networks [38, 5]. As the next step, incorporating new material as the robot body can be achieved by including material from the environment, e.g., by combining soil and probe-deposited linker for root-inspired robots [34]. Functional body components such as grippers or sensing arrays could be retrofitted via physical additions (upgrades) on the robot body surface with advanced 3D printing [19]. In fluidic devices, tactile sensing could be retrofitted by measuring the single fluidic input [45]. All in all, current strategies to extend the robot’s capabilities during deployment rely on existing hardware or tailored retrofitting, with no new structurally and functionally intertwined hardware generation.
@@
When creating new hardware, the main bottleneck is the material - suitable material compositions should be placed at the optimal location. In nature, material transport in the organism is solved with passive mechanisms (diffusion) to reach molecule-level interactions (e.g., receptor binding) and active mechanisms (advection, particularly circulation) to minimize diffusion distance [25]. In robotics, the open circulatory systems of simpler organisms like arthropods serve as architectural models for designing material transport networks. They consist of a pumping element (heart), flow-directing arteries and hemolymph as a circulatory fluid that bathes the internal organs and infuses wings, as in Lepidoptera (Figure 1a). Insect wings are living structures [35], in which the hemolymph supports exoskeleton/cuticle hydration for stiffness modulation [26], thermal regulation [39], aerodynamic properties [26, 8] and chemical[14] and mechanical[29] signalling, providing an accessible (with effectively 2D fluid flow) case study for vascularised robotics.
@@
Artificial interpretations of vascularisation are highlighted by progress in microfluidics, e.g, organ-on-a-chip or drug testing solutions [4], as well as structures for liquid transport [13]) and patterning [23]. In robotic systems, a multifunctional internal liquid was circulated throughout a lionfish-inspired robot, simultaneously powering and actuating it [3]. In terms of intra-fluid communication, Garrad et al demonstrated injecting patterned sequences of conducting and insulating fluid into a soft tube produces measurable resistance changes across embedded electrodes, encoding the fluidic input as a binary electrical signal [17]. In our previous work [40], we showed the integrative role of internal liquid for simultaneous body compliance modulation and muscle activation while reaching protovascular level, highlighting the capability of internal liquid to enable physically intelligent robotic systems. For living mammalian cells, oxygen diffusion distance is in the range of to [32] which gives clear motivation to develop vascular structures for cell viability in biohybrid systems [15] and hints to potential architectures for material-level in-situ morphogenesis. Robotic vasculature (a.k.a. embedded functional vias [6]) are proposed to alter robots’ body properties in response to environmental conditions, offering a pathway to mimic phenotypic plasticity in organisms [6]. However, there are currently no reports on robotic vasculatures that close the loop from an environmental stimulus to a bodily adaptation, in part due to fabrication (printing) challenges.
@@
Here, we report the ex novo generation of functional hardware components through an integrated vascular–materials strategy. To enable controlled matter transport, we introduce a vascularization approach that couples long‑range advective flow with localized infusion. At the component level, we demonstrate that polypyrrole (PPy) synthesized via environmental UV‑induced photopolymerization is a feasible UV‑responsive receptor material. We validate our framework on a moth‑inspired (Figure 1b, left) flapping‑wing robotic platform (Figure 1b, right), where a circulatory fluid is routed through an engineered vascular network and used to deliver PPy precursors to the receptor site to be photopolymerized. This work advances autonomous robotic function in complex environments by extending morphological computation to the molecular scale, and provides a materials‑centric and intrinsic dynamics driven perspective on neurovascular integration in robotic systems.
@@ 2.1 System Design
Figure 1c outlines the key steps of ex novo component generation using receptogenesis as a case study. We create functional components in situ by circulating precursors through an internal artificial vasculature and photopolymerising them with external stimulus (UV light). Inspired by the open circulatory systems of Lepidoptera (Figure 1b), we implement a 3D printed artificial vascular network that distributes material precursors throughout the body via global advection and local fluid infusion. We use polypyrrole (PPy) precursors for UV-sensitive receptor photopolymerization. This newly formed component functions as a light‑responsive receptor capable of triggering downstream behaviours such as LED indication or wing flapping (Figure 1d). In our system, the stimulus that initiates receptogenesis and the signal being detected are incidentally the same, though the approach generalizes to a much broader class of stimuli and signals.
We chose polyethylene terephtalate glycol (PETG) as the 3D printing polymer as it is commercially available and transmits UV-light starting from [2]. Pyrrole (Py) as an organic solvent infuses PETG and can be photopolymerised into PPy [31], during which the transparent Py forms dark UV-sensitive [24] PPy clusters. The perceived colour change is a convenient way of observing the synthesis process.
@@
In our work, precursors are a liquid mix of pyrrole (acting as both monomer and solvent), a bis(4-methylphenyl)iodonium hexafluorophosphate initiator, and cellulose acetate propionate (CAP) as a structural modulator. We conceptualize this chemical inventory as a dual-source resource: it may be integrated as an on-board reservoir to ensure operational readiness, or alternatively, sequestered from the ambient environment via harvesting mechanisms. While the logistics of precursor procurement fall outside the primary scope of this study, the capacity for both intrinsic storage and extrinsic acquisition presents a robust framework for autonomous system evolution.
@@
Surface-level modifications (”add-ons”) often act as functional obstructions (analogous to intravascular clots) - limiting the functional versatility of the vascular system. Instead, we focus on the physical update (chemical synthesis) in the PETG matrix as opposed to the immediate vascular surface. In macroscale robots, however, diffusion alone is too slow to deliver precursor uniformly throughout the volume, limiting the rate at which the material can fully activate. To bypass this bottleneck, we implement advection that ensures rapid, uniform activation. We developed a global logistics network combining rapid fluid circulation through a vascular scaffold complemented by slower, diffusion-limited local precursors delivery via infusion into the PETG structure(Figure 2a). Fused deposition modelling (FDM) printing of PETG produced logistic highways (veins) by explicit definition and a distributed permeability structure (microchannels) by designed infill pattern (Figure 2b, Supplementary Figure S1). In the microchannels (Figure 2c), the circulatory fluid infused the volume of the PETG scaffold (Figure 2d), enabled by the large surface area in contact with the fluid. Pyrrole (Py) solution infused the PETG (exposure time approximately ) to an infusion depth of approximately as seen from the element contrast in the SEM image in Figure 4d (bottom right). The formed Py-PETG layer was approximately 20% of the print line’s thickness ( radius). This enables us to achieve connectedness across multiple orders of scale (tens of centimeters via global transport and micrometers via infusion) using a single fabrication method.
@@ 2.2 Receptogenesis by In-vasculature Molecular Assembly
Receptogenesis is achieved via photopolymerization of Py precursors into a UV-receptive area induced by environmental UV-stimulus (Figure 3a). The circulatory fluid of precursors infuses the PETG structure (Figure 3b, Supplementary Movie 1) and forms an infused PETG (Py-PETG) region in the structure. Photopolymerization of Py is self-inhibiting as the synthesis takes place at the intersection of circulatory fluid infusion front and UV-light. Since PPy and the precursors absorb UV light, synthesis is preferentially localized at UV-exposed surfaces. Visually the synthesised area is dark colored. To characterise it, we monitored the absorption in-situ during synthesis (Figure 3c). The absorption rate was and irreversible. We expect that the PPy formed clusters in the Py-PETG matrix that self-inhibit further growth when the immediate vicinity has been depleted of precursors as previously discussed in literature [31]. This is consistent with non-zero final transmittance of in Figure 3c. Conduction combines electronic conduction of PPy clusters and electrolytic conduction of initiator in the PETG matrix surrounding the clusters, similarly to described elsewhere [43]. We demonstrated UV-PPy photolithography using a PETG negative contact mask (Figure 3d), achieving resolution , limited by optical scattering by surface roughness.
@@
PPy is sensitive to UV irradiation[16], which can photo-excite charge carriers and modify the population or dynamics of polaron or bipolaron states (Figure 3e), leading to impedance decrease in our PPy-based system. The impedance change was registered by an embedded microprocessor unit featuring a custom bipolar pulse technique combined with voltage division by a constant external resistor (Figure 3f, left). The lower transient voltage across the receptive area (PPy-PETG), versus Py-PETG, showed higher conductance, whereas the lower voltage rate captured an increased capacitance (Figure 3f, right). Further, we encoded receptor response as a differential digital signal with LED readout, given the transient nature of polarons: the sensor response is characterized by its rate of change rather than absolute value. This parallels biological receptors, as we are not interested in absolute (impedance) values, but rather the change. As seen in Figure 3g, a drop in impedance corresponds to UV irradiation that was independently visualized with a red LED blinking.

### 19 May 2026 v2

@@ Abstract
Equipping robotic systems with the capacity to generate ex novo hardware during operation extends control of physical adaptability. Unlike modular systems that rely on discrete component integration pre- or post-deployment, we envision the possibility that physical adaptation and development emerge from dynamic material restructuring to shape the body’s intrinsic functions. Drawing inspiration from circulatory systems that redistribute mass and function in biological organisms, we utilize fluidics to restructure the material interface, a capability currently unpaired in robotics. Here, we realize this synthetic growth capability through a vascularized robotic composite designed for programmable material synthesis, demonstrated via receptogenesis - the on-demand construction of sensors from internal fluid reserves based on environmental cues. By coordinating the fluidic transport of precursors with external localized UV irradiation, we drive an in situ photopolymerization that chemically reconstructs the vasculature from the inside out. This reaction converts precursors with photolatent initiator into a solid dispersion of UV-sensitive polypyrrole in PETG, establishing a sensing modality validated by a characteristic decrease in electrical impedance. The newly synthesized sensor closed a local control loop to regulate wing flapping in a moth-inspired robotic demonstrator. This physical update increased the robot’s capability in real time. Material-level functional restructuring of the vascularized robot body provides a proof-of-concept materials basis for ex novo hardware generation in situated robotic systems - a step toward situated robots in which a reaction to environmental stimuli autonomously produces hardware updates to match new environmental demands.
@@ 1 Introduction
In unstructured environments, critical physical signals are not static features that can be precisely registered by a dedicated sensor, but rather emergent phenomena arising from the robot’s unique interactions with the environment [32, 1]. We need more focus on situated robots [48] that are deeply intertwined with their immediate surroundings and leverage the environment itself to guide their behaviour and expand their physical capabilities. In situ generation of new hardware within the robot body based on environmental cues would couple environmental information with physical updates, adaptively creating structures that transduce situation-specific signals emerging from the environment.
Prior approaches to physical adaptation during deployment differ in the extent of bodily changes. Some rearrange existing hardware — through compliance and deformation [39], appendage repurposing for locomotion plasticity [44], shape-morphing limbs [7] or body-wide morphological reconfiguration [33] for multi-environment transitions, or software remapping of existing sensors, e.g., repurposing navigational accelerometers for gravimetry in the Curiosity rover [26]. Self-healing systems restore the body after damage, knitting severed material back together through reversible bonds in embedded polymer networks [45, 5]; the endpoint, however, is a recovered prior state rather than a new function. Growth-based robots incorporate genuinely new matter, as in root-inspired systems that bind ambient soil with a tip-deposited linker [40]; here, material is accreted at an external growth front rather than synthesized within the existing body. Retrofitting strategies place new functional components onto the robot via in-situ 3D printing [21], or repurpose existing fluidic channels to extract sensory information from a single pressure input [53]; in both cases, the added capacity remains structurally discrete from the body it extends, acting as an add-on rather than an integral part that invokes new behavior. Across all these approaches, the new capacity is either a rearrangement of what was already there, a return to the prior state, or an external appendage — none synthesize hardware that is structurally and functionally continuous with the body, formed internally, and providing a new qualitative function.
@@
Synthesizing hardware internally to the body shifts the problem from component placement to coupled transport and assembly: precursors must travel centimeters through a vascular network, then convert into function across micrometres through molecular assembly. In nature, analogous material transport issues are solved with passive mechanisms (diffusion) for molecule-level interactions such as receptor binding, and active mechanisms (advection, particularly circulation) to minimize diffusion distance [29]. In robotics, the open circulatory systems of simpler organisms such as arthropods serve as architectural models for material transport networks. They consist of a pumping element (heart), flow-directing arteries, and hemolymph — a circulatory fluid that bathes the internal organs and infuses the wings, as in Lepidoptera (Figure 1a). Insect wings are living structures [41], in which hemolymph supports cuticle hydration for stiffness modulation [31], thermal regulation [46], aerodynamic properties [31, 8], and chemical [14] and mechanical [34] signalling, providing — with effectively 2D fluid flow — an accessible case study for vascularised robotics.
@@
Artificial interpretations of vascularization are highlighted by progress in microfluidics, e.g., organ-on-a-chip or drug testing solutions [4], as well as structures for liquid transport [13] and patterning [27]. In robotic systems, a multifunctional internal liquid was circulated throughout a lionfish-inspired robot, simultaneously powering and actuating it [3]. In terms of intra-fluid communication, Garrad et al demonstrated that injecting patterned sequences of conducting and insulating fluid into a soft tube produces resistance changes across embedded electrodes, encoding the fluidic input as a binary electrical signal [18]. In our previous work [47], we demonstrated the integrative role of internal liquid in simultaneously modulating body compliance and activating muscles, reaching protovascular levels, highlighting its enabling role in physically intelligent robotic systems. For living mammalian cells, oxygen diffusion distance is in the range of to [38], which gives clear motivation to develop vascular structures for cell viability in biohybrid systems [15] and hints to potential architectures for material-level in-situ morphogenesis. Robotic vasculature (also termed embedded functional vias [6]) is proposed to alter robots’ body properties in response to environmental conditions, offering a pathway to mimic phenotypic plasticity in organisms [6]. However, no robotic vasculature has yet realized this plasticity in a local closed loop: from an environmental stimulus, through material change, to a new functional capability of the body.
Taken together, the existing strategies for robots’ physical adaptability and vascular architectures have remained disconnected. Vascularization enables through-embodiment augmentation: new functional material is synthesized within the existing body structure, resulting in ex novo hardware capability that the robot did not possess at manufacture, without altering the perimeter or compromising structural integrity. When the ex novo component is a sensory receptor, it grants the robot access to environmental signals it could not previously detect; we term this receptogenesis. Both terms (ex novo hardware generation and receptogenesis) describe robots’ operational capability, allowing the vast chemical synthesis toolbox to functionally augment robots.
@@
Here, we report the ex novo hardware generation through an integrated vascular–materials strategy, realized as receptogenesis. We introduce a two-scale vascularization strategy in which a single fabrication method bridges global and local transport: 3D-printed veins carry precursors across the decimeter-scale body, while infusion drives them into the surrounding PETG matrix down to . At the component level, we show that polypyrrole (PPy) synthesized via UV-induced photopolymerization is a feasible UV-responsive receptor material. We validate our framework on a moth‑inspired (Figure 1b, left) flapping‑wing robotic platform (Figure 1b, right), where a circulatory fluid is routed through an engineered vascular network and used to deliver PPy precursors to the receptor site to be photopolymerized. The newly synthesized receptor closes a control loop regulating wing flapping and visual signaling (Figure 1d), demonstrating a functional capability the robot did not possess prior to deployment. This work advances autonomous robotic function in complex environments by extending morphological computation to the molecular scale, and provides a materials‑centric and intrinsic dynamics driven perspective on neurovascular integration in robotic systems.
@@ 2.1 System Design
Figure 1c outlines the key steps of ex novo component generation using receptogenesis as a case study. We create functional components in situ by circulating precursors through an internal artificial vasculature and photopolymerising them with external stimulus (UV light). Inspired by the open circulatory systems of Lepidoptera (Figure 1b), we implement a 3D printed artificial vascular network that distributes material precursors throughout the body via global advection and local fluid infusion. We use polypyrrole (PPy) precursors for UV-sensitive receptor photopolymerization. This newly formed component functions as a light‑responsive receptor capable of triggering downstream behaviours such as LED indication or wing flapping (Figure 1d). In our system, the stimulus that initiates receptogenesis and the signal being detected are incidentally the same, though the approach generalizes to a much broader class of stimuli and signals. Even when narrowly considering PPy, the case study material in this paper, the synthesis result is based on the precursor choice and environmental (synthesis) conditions, leading to a wide range of possible outcomes (e.g., towards bioelectronic interfacing [11], ionic sensitivity [49], corrosion resistance [10], or electromagnetic wave shielding [24]). Even considering the light-patterning technique, external stimulus can modulate photocrosslinking patterns for microphase separation of a copolymer, achieving stiffness patterning in a chemically identical material [35], providing possible application scenarios in our vascularized system.
This work focuses on the ’small’ loop — local, stimulus-mediated synthesis in a vascular system primed to encounter the stimulus. The ’large’ loop (central decision-making/autonomy, e.g., priming the body with task-relevant precursors, downstream behavior) is implicit: in the experiments reported here, priming was performed via user input, and the body was then exposed to controlled stimuli that activate the primed receptor site. Within the small loop, no central involvement is required during execution — body articulation, self-occlusions, surface orientation, and ambient light gradients would collectively determine where and when polymerization occurs in a deployed system, a role played in our demonstration by a physical mask. Synthesis is therefore shaped by the conditions of each encounter, intrinsically coupling the generated hardware to the environment in which it operates.
@@
In our work, precursors are a liquid mixture of pyrrole (Py, acting as both monomer and solvent), a bis(4-methylphenyl)iodonium hexafluorophosphate initiator, and cellulose acetate propionate (CAP) as a structural modulator. We conceptualize this chemical inventory as a dual-source resource: it may be integrated as an on-board reservoir to ensure operational readiness, or alternatively, sequestered from the ambient environment via harvesting mechanisms. While the logistics of precursor procurement fall outside the primary scope of this study, the capacity for both intrinsic storage and extrinsic acquisition presents vascularization as a robust framework for autonomous system evolution. We chose polyethylene terephtalate glycol (PETG) as the 3D printing polymer as it is commercially available and transmits UV light starting at [2]. Pyrrole as an organic solvent infuses PETG and can be photopolymerised into PPy [36], during which the transparent Py forms dark UV-sensitive [28] PPy clusters. The perceived color change evidences the synthesis process.
@@
We focus on the physical update (chemical synthesis) in the PETG matrix, opposed to surface-level modifications (”add-ons”) that often act as functional obstructions (analogous to intravascular clots), limiting the functional versatility of the vascular system. In macroscale robots, diffusion alone is too slow to deliver precursors uniformly throughout the volume, limiting the full activation rate. To bypass this bottleneck, we implement advection that ensures rapid, uniform activation. We developed a global logistics network combining rapid fluid circulation through a vascular scaffold complemented by slower, diffusion-limited local precursors delivery via infusion into the PETG structure (Figure 2a). Fused deposition modelling (FDM) printing of PETG produced logistic highways (veins) by explicit definition and a distributed permeability structure (microchannels) by designed infill pattern (Figure 2b, Supplementary Figure S1, Supplementary Movie 1). In the microchannels (Figure 2c), the circulatory fluid infused the PETG scaffold (Figure 2d), enabled by the large surface area in contact with the fluid. Precursor mixture infused the PETG (exposure time approximately ) to an infusion depth of approximately as seen from the element contrast in the SEM image in Figure 4d (bottom right). The formed Py-PETG layer was approximately 20% of the print line’s thickness ( radius). This enables us to achieve connectedness across multiple orders of scale (tens of centimeters via global transport and micrometers via infusion) using a single fabrication method.
@@ 2.2 Receptogenesis by In-vasculature Molecular Assembly
Receptogenesis is achieved via photopolymerization of Py precursors into a UV-receptive area induced by initiator decomposition under UV-stimulus (Figure 3a). The circulatory fluid of precursors infused the PETG structure (Figure 3b, Supplementary Movie 2) and formed an infused PETG (Py-PETG) region in the structure. Py-PETG formed a solid solution, in which the precursor molecules retained sufficient mobility to reorganize within the matrix, enabling residual polymerization after the initial UV exposure (synthesis event). Since PPy and the precursors absorb UV light, synthesis was localized at UV-exposed surfaces. The reaction area progressively and irreversibly darkened at an absorption rate of , monitored at in-situ during synthesis (Figure 3c). Based on prior literature of photopolymerization [36, 22] and known PPy morphology from other chemical synthesis methods [23, 30], PPy most likely formed clusters in the Py-PETG matrix that self-inhibit further growth when the immediate vicinity has been depleted of precursors. This is consistent with non-zero final transmittance of in Figure 3c. Conduction combines electronic conduction of PPy clusters and electrolytic conduction of initiator in the PETG matrix surrounding the clusters, forming unique formations each time, in agreement with concept described elsewhere [51]. Absolute resistance varied widely between samples (), consistent with sample-specific PPy microstructure, whereas relative UV-sensitivity remained consistent (detailed below). We therefore report UV-responsiveness as a relative resistance change rather than an absolute value. We demonstrated UV-PPy photolithography using a PETG negative contact mask (Figure 3d) to show that UV-induced synthesis was spatially localized and did not propagate beyond the exposed area, achieving resolution , limited by optical scattering by surface roughness.
@@
As the vicinity of the receptive area retained abundant Py monomers after self-inhibiting PPy photosynthesis, the Py monomers reduced the newly synthesized PPy into an electrically nonconductive form. PPy is sensitive to UV irradiation [16]: the sensing characteristics of the receptive area (PPy-PETG) derive from the transient formation of (bi)polarons: UV exposure photo-excites PPy by separating charge carriers and modifying the population or dynamics of polaron or bipolaron states (Figure 3e). The induced positive charge shifts PPy into its oxidized, electrically conductive form, leading to resistance decrease in our PPy-based system. The impedance drop was proportional to the LED pulse (a proxy for UV dose, Figure 3f), indicating that (bi)polaron population was proportional to the incident UV dose. The conductive state is transient, decaying through two routes: charge consumed by further polymerization of the abundant surrounding Py (extending the existing PPy chain), and carrier recombination limited by charge mobility along the chain (timescale ). Over repeated cycling, both sensitivity (Figure 3g) and baseline resistance values increased, consistent with residual polymerization of Py: additional PPy generates more (bi)polarons, increasing sensitivity, whereas a decreased amount of mobile charge carriers increases system resistance. This constant evolution of the sensor is expected and embodies the desired element of experience to the situated sensor, different from the expectedly constant performance of conventional (industrial) sensors. The system retained functionality after more than four months of storage, demonstrating operational stability on this timescale. For the receptor, the impedance change of the receptive area was registered by an embedded microprocessor unit featuring a custom bipolar pulse technique combined with voltage division by a constant external resistor (Supplementary Figure S2). Further, we encoded receptor response as a differential digital signal with LED readout, given the transient nature of polarons: the receptor response is characterized by its rate of change rather than absolute value, as described above. This parallels biological receptors, as we are not interested in absolute (impedance) values, but rather the change. As seen in Figure 3h, a drop in impedance corresponds to UV irradiation that was independently visualized with a red LED blinking.